An income property has 5000 square feet of leasable space, and rents for $8.00 per square foot per year. Vacancy and credit loss are estimated at 4% and operating expenses at 38% of potential gross income. Lenders require a 1.2 Debt Coverage Ratio. Mortgage is available at terms of 25 years, at 15%, payable annually. The investor requires a 14% rate of return on the projects before tax equity cash flow.

If other lenders required a Debt Coverage Ratio of 1.4, how would the maximum offering price be affected?

  1. It would go up.
  2. It would go down.
  3. It would remain the same.
  4. It would not be relevant to the price.
Answer #1

Given,

Leasable area= 5000 Sq. ft.

Rent per sq. foot= $8

Therefore, potential gross income= 5000*8 = $40,000

Also given, vacancy and credit loss= 4% and operating expense= 38% of gross income

Therefore, net operating income (NOI)= 40,000*(1-42%)= $23,200

Also given, Debt Coverage Ratio (DCR) Required=1.2

Therefore, maximum yearly payment= NOI/DCR= 23,200/1.2= $19,333.33

Net yearly cash flow= NOI-Yearly payment= $23,200-$19,333.3333 = $3,866.6667

Mortgage term: Period=25 years, Interest rate=15%

Required return on equity= 14%

Down payment corresponding to net cash flow= $3,866.6667*PVIFA(14%,25)

=3866.6667* 6.873 = $26,575.60

Loan amount corresponding to annual payment of $19,333.3333 = 19,333.3333*PVIFA(5%,25)

=19,333.3333*6.464 = $124,970.67

Maximum offering price= Loan + Down payment= $124,970.67 + $26,575.60 = $151,546.27

If the Debt Service Ratio =1.4,

Therefore, maximum yearly payment= NOI/DCR= 23,200/1.4= $ 16,571.43

Net yearly cash flow= NOI-Yearly payment= $23,200-$ 16571.43 = $ 6,628.57

Mortgage term: Period=25 years, Interest rate=15%

Required return on equity= 14%

Down payment corresponding to net cash flow= $ 6,628.57*PVIFA(14%,25)

= 6628.57* 6.873 = $ 45,558.17

Loan amount corresponding to annual payment of $ 16,571.43= 16,571.43*PVIFA(5%,25)

=16,571.43*6.464 = $107,117.71

Maximum offering price= Loan + Down payment= $107,117.71 + $ 45,558.17 = $152,675.89

As seen above, if Debt Service Ratio is 1.4, maximum offering price will go up. The answer is option (a).